# Minchul Kim, Anil Jain, Xiaoming Liu

### v2

@@ 4 Advances in Deep Face Recognition
Over the past decade, deep face recognition has experienced remarkable progress, driven primarily by three key factors: the development of loss functions, the availability of large-scale and diverse datasets, and advances in neural network architectures. Together, these innovations have dramatically improved the ability of models to learn highly discriminative, robust facial representations.
@@ 4.1 Loss Functions
The choice of loss function is critical in training deep face recognition models, as it directly guides the network to learn discriminative feature embeddings suitable for distinguishing between a vast number of identities. Among the key innovations that drove the advancement in FR, the highest number of publications have come from the advances in the loss function. Several distinct paradigms for loss function design have emerged.
Contrastive Learning Approaches: One major family of loss functions employs contrastive learning principles, directly shaping the embedding space by optimizing relative distances between samples. The seminal FaceNet [14] introduces the triplet loss, designed to ensure that an anchor sample’s embedding is closer to its positive (same identity) counterpart than to any negative (different identity) sample by a predefined margin, typically in the Euclidean space.
While effective, triplet loss faces challenges in sampling informative triplets. Many randomly chosen triplets provide weak gradient signal, making training inefficient or necessitating complex hard-negative mining strategies. To address this, proxy-based methods are proposed. Techniques like Proxy Anchor Loss [57], originating from general deep metric learning, associate each class with learnable proxies (representative vectors), simplifying the loss computation by comparing samples to these proxies rather than exhaustively searching for pairs or triplets within a batch.
@@
DeepID2+ [58] 2014 Joint ID + verification loss for robust features Complex training with dual supervision losses
@@
CenterFace [59] 2016 Center loss enhances intra-class compactness No explicit inter-class separation; needs tuning
@@
SphereFace [18] 2017 Angular margin enforces hyperspherical separation Training instability from angular multiplicity
@@
L2-Face [60] 2017 L2 norm constraint improves angular discrimination Needs careful radius tuning; no margin enforcement
@@
ArcFace [20] 2018 Additive angular margin boosts inter-class separation Fixed margin may hurt low-quality samples
@@
CosFace [19] 2018 Cosine margin improves class separability stably Uniform margin not adaptive; needs tuning
@@
SeqFace [61] 2018 Sequence-aware loss improves temporal supervision Needs sequence data; dual loss increases complexity
@@
Git Loss [62] 2018 Unified softmax + center loss boosts discrimination Extra tuning and complexity with marginal gain
@@
MagFace [63] 2021 Feature norm models quality for adaptive margin Complex loss and quality-norm assumptions
@@
AdaFace [22] 2022 Dynamic margin based on feature norm quality Relies on norm-quality link and tuning
@@
ElasticFace [64] 2022 Elastic margin adapts to feature variability Stochastic margins add tuning and training cost
@@
UniFace [65] 2023 Similarity threshold improves verification alignment Global constraints increase optimization cost
@@
UniTSFace [66] 2023 Sample-to-sample loss optimizes verification Pairwise loss and threshold learning cause overhead
@@
UCFace [67] 2024 Uncertainty and probability density aware contrastive learning Cannot be used by itself, must be accompanied by margin loss
@@
LAFS [68] 2024 Landmark based SSL pretraining helps face recognition Loss depends on pretrained model and the landmark quality.
@@
Further refining the contrastive approach, Supervised Contrastive Learning (SupCon) [69] generalizes the loss to leverage all positive samples available for an anchor within a batch, contrasting them against all negative samples. This more data-efficient approach has been successfully applied to FR, for instance, in UCFace [67]. Other works adapt contrastive ideas for specific goals: Open-Set Biometrics [70] focuses on improving open-set performance by explicitly minimizing scores between non-mated pairs, while CAFace [71] uses a contrastive-style cosine similarity loss to enforce consistency between embeddings of low-quality images and their high-quality counterparts, promoting quality invariance. Another work in this category, Related efforts also explore optimizing embedding spaces to better align with recognition objectives, such as in [72], where features are learned to directly improve performance metrics for face identification and verification separately.
@@
Margin-based Softmax Losses: A dominant and highly successful approach in deep FR involves modifying the standard softmax cross-entropy loss to directly enhance feature discriminability. The core motivation is to learn embeddings that exhibit smaller intra-class variations (features of the same person are close together) while simultaneously maximizing inter-class separation (features of different people are far apart).
@@
$\mathcal{L}_{CE}(\mathbf{x}_{i})=-\log\frac{\exp(\mathbf{W}_{y_{i}}^{T}\mathbf{z}_{i}+b_{y_{i}})}{\sum_{j=1}^{C}\exp(\mathbf{W}_{j}^{T}\mathbf{z}_{j}+b_{j})},$ (1)
@@
where $\mathbf{W}_{j}$ is the weight vector for the $j$-th class, $b_{j}$ is the bias term, and $C$ is the total number of classes or identities in the training set. While effective for classification, this formulation doesn’t explicitly enforce the metric learning objective crucial for face recognition where we encounter identities not seen during training.
@@
An early work moving in this direction is Center Loss [59], which adds an auxiliary loss term to the standard softmax. This term penalizes the Euclidean distances between the deep features and their corresponding learned class centers, directly encouraging intra-class compactness. A significant breakthrough comes with the normalization of both feature embeddings ($\|\mathbf{z}_{i}\|=1$) and classification weights ($\|\mathbf{W}_{j}\|=1$, and setting $b_{j}=0$). This reformulation, pioneered by SphereFace [18], maps the optimization problem onto a hypersphere where the dot product $\mathbf{W}_{j}^{T}\mathbf{z}_{i}$ becomes equivalent to $\cos\theta_{j}$, the cosine of the angle between the feature vector $\mathbf{z}_{i}$ and the weight vector $\mathbf{W}_{j}$. A scaling factor $s$ is typically introduced to control the radius of the hyperspherical feature space. The loss then becomes:
$\mathcal{L}_{cos}(\mathbf{x}_{i})=-\log\frac{\exp(s\cdot\cos\theta_{y_{i}})}{\sum_{j=1}^{C}\exp(s\cdot\cos\theta_{j})}.$ (2)
@@
Building on this normalized angular space, the key innovation is the introduction of explicit margins to make the learning objective more stringent. CosFace [19] introduces an additive cosine margin ($m$) by modifying the target logit to $s\cdot(\cos\theta_{y_{i}}-m)$. ArcFace [20] proposes an additive angular margin ($m$) by modifying the target angle itself, resulting in a target logit of $s\cdot\cos(\theta_{y_{i}}+m)$. Both approaches effectively create a decision boundary gap, forcing learned features for the same identity to cluster more tightly in the angular space, thereby significantly improving discriminative power. A visual comparison of contrastive triplet loss and margin-based normalized softmax loss is illustrated in Fig. 8, highlighting how each paradigm optimizes the embedding space to enhance FR performance.
@@
Subsequent research further refines these margin-based concepts. MagFace [63] proposes leveraging the magnitude of the feature vector (before normalization) as an indicator of face image quality, incorporating an auxiliary loss to promote larger magnitudes for higher-quality samples. AdaFace [22] addresses the challenge posed by low-quality or difficult samples by introducing an adaptive margin function. It dynamically adjusts the margin stringency based on image quality indicators, reducing the negative impact of potentially unrecognizable faces in the training process.
@@
These advancements in margin-based softmax losses lead to remarkable performance gains, pushing verification accuracy on high-quality benchmarks like LFW [26] and CFP-FP [27] towards saturation (often exceeding 99%). This success shifts the community’s focus towards improving performance in more challenging, unconstrained scenarios, particularly those involving low-quality images, as represented by benchmarks like IJB-S [25, 73].
@@
Further refinements continued to explore margin dynamics; for example, ElasticFace [64] introduces randomized margins for greater flexibility during training, while UniFace [65] proposes the Unified Cross-Entropy (UCE) loss specifically aiming to guarantee a clear separation threshold between positive and negative pairs.
@@
Margin-based softmax variants [20, 22, 65] currently dominate SoTA results. The specific choice among these often depends on the dataset characteristics (e.g., presence of low-quality images) and desired properties. Contrastive methods remain a helpful auxiliary loss, on top of margin-based softmax losses. A summary of various loss functions are shown in Tab. I.
@@
A significant effort has focused on improving model interpretability, e.g., understand how the network makes decisions. Towards this, Yin et al. [74] propose spatial and feature activation diversity losses. These encourage the network to learn more structured representations where different spatial activations may correspond to different facial aspects, while also making these interpretable features discriminative and robust to occlusions. Similarly, the Explainable Channel Loss (ECLoss), also framed as Activation Template Matching Loss [75], encourages specific channels within convolutional layers to specialize in detecting distinct facial parts (e.g., eyes) without explicit part annotations, thereby providing a direct interpretation of channel function.
Knowledge distillation (KD) offers another avenue, often targeting specific challenges. For low-resolution FR, Attention Similarity KD (A-SKD) [76] transfers teacher attention maps to guide the student’s focus. For efficient large-scale training, Li et al. [77] develop feature-based KD techniques, like reverse distillation, that importantly remove the need for identity supervision for the student, saving memory while addressing the teacher-student ’intrinsic gap’. These examples highlight the use of specialized loss functions and KD strategies to imbue models with desirable properties like explainability, robustness, or training efficiency, complementing the primary task of recognition.
@@ 4.2 Datasets
The availability and scale of training data have been pivotal factors driving the remarkable progress in deep learning [15]. Publicly available large-scale face datasets (e.g., MS-Celeb-1M, VGGFace2) spurred rapid advancements in the mid-2010s. Tab. II provides an overview of several influential datasets commonly used in the field, detailing the number of images and unique identities they contain. A clear trend emerges from this summary: a dramatic increase in dataset size over time. Early benchmark datasets like CASIA-WebFace [53] offer around half a million images from ten thousand subjects. In contrast, subsequent collections such as MS1MV2/V3 [78], Glint360K [79], and particularly the WebFace series [55], have pushed these numbers significantly higher, culminating in WebFace42M with over 42 million face images spanning 2 million identities. This growth reflects the community’s understanding that larger and more diverse datasets are crucial for training accurate and generalizable FR models. It is important to note that MS-Celeb-1M and VGGFace2 have been redacted from the dataset creators due to lack of user consent.
Figure 10: Plots showing the growth of face recognition datasets over time. The left plot illustrates the number of images (in millions), and the right plot shows the number of subjects (in thousands) for each dataset.
@@
It is important to note the origin and labeling methodology of many of these large-scale datasets. A significant portion, including prominent datasets like MS-Celeb-1M and the WebFace series, are curated by collecting images from publicly accessible sources on the Internet, often leveraging search engines or social media platforms. Consequently, the identity labels associated with these images are frequently “pseudo-labels,” because web searches of celebrities may return different subject images. Due to the volumn of dataset, the labels are generated through automated clustering algorithms or matching, rather than manual verification. While efforts are made to clean and refine these labels, noise and inaccuracies can persist. Some approaches, like that used for WebFace260M [55], employ iterative self-labeling and retraining of specialized labeler models to improve the quality of these pseudo-labels over multiple cycles. Since benchmark datasets [26, 27, 29, 23, 24] are also curated from public web, training datsets need to ensure that the identities in training and test sets do not overlap.
@@
Also, the practice of web-scraping facial images raises significant privacy and ethical concerns within the research community and society at large [80], as individuals may not have provided explicit consent for their images to be used in this manner for developing and training recognition systems. This issue remains an active area of discussion and necessitates careful consideration of data governance and ethical guidelines moving forward.
@@
In addition to the large-scale 2D image datasets collected primarily from the web, the field also utilizes 3D face datasets [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92]. These datasets capture the geometric structure of the face, often along with texture information, using specialized acquisition techniques like 3D scanners, structured light, or multi-camera stereo systems. A key distinction is that 3D datasets are typically collected under controlled laboratory settings with the explicit consent of the participants. This controlled acquisition allows for high-quality, precise capture of facial shape in RGBd (depth), which can offer inherent robustness advantages against variations in pose and illumination compared to 2D images.
@@
3D face recognition has also emerged as a parallel research track, with dedicated benchmarks and evaluation campaigns [93]. Notably, datasets such as FRGC [93] and Lock3DFace [94], which leverages low-cost RGB-D sensors like Kinect, have been widely used in the community to advance recognition algorithms under realistic conditions.
@@
However, the process of 3D data acquisition is significantly more complex, time-consuming, and expensive. Consequently, the volume of available 3D face data, both in terms of the number of scans and the number of unique subjects, is substantially smaller compared to the massive scale achieved by web-scraped 2D datasets. This difference in scale limits the use of 3D face datasets for training the deep models that dominate current FR research, although they remain valuable for specific research tasks, evaluation, and applications where 3D information is critical. For general face recognition applications, such as in law enforcement, immigration, or airport screening, RGB cameras offer a more practical solution considering legacy databases and return on investment.
TABLE II: Overview of face recognition datasets, used primarily for training deep networks, with corresponding image and subject counts, year of release, and authentication performance reported as TAR@FAR=0.01%.
Dataset Year # Images # Subjects Max IJB-C TAR@FAR=0.01%
CASIA-WebFace [53] 2014 0.5M 10K 89.84 (UniTSFace)
VGGFace2 [54] 2018 2.6M 2.6K
MS1MV2 [78] 2019 5.8M 85K 95.67 (AdaFace)
MS1MV3 [78] 2019 5.1M 93K 95.10 (ArcFace)
Glint360K [79] 2021 17M 360K 97.20 (PartialFC)
WebFace4M [55] 2021 4.2M 200K 96.03 (AdaFace)
WebFace12M [55] 2021 12M 600K 97.82 (ViT KP-RPE)
WebFace42M [55] 2021 42M 2M 97.99 (UniTSFace)

### v3

@@ 4 Advances in Deep Face Recognition
Deep face recognition has advanced through improved loss functions, large diverse datasets, and better neural architectures, enabling more discriminative representations.
@@ 4.1 Loss Functions
Before deep learning, FR relied on analytical methods lacking desired discriminative power. Eigenfaces [turk1991face] used Principal Component Analysis to maximize variance, and Fisherfaces [belhumeur1997eigenfaces] applied Linear Discriminant Analysis to enhance class separability, both requiring costly covariance matrix diagonalization [martinez2001pca]. Deep learning replaced these with gradient based optimization and specialized loss functions [wang2017normface, liu2017sphereface, wang2018cosface, deng2019arcface, huang2020curricularface, kim2022adaface] that directly enforce inter-class separability and intra-class compactness.
The choice of loss function is critical in training deep FR models, as it directly guides the network to learn discriminative feature embeddings suitable for distinguishing between a vast number of identities. Among the key innovations that drove the advancement in FR, the highest number of publications have come from the advances in the loss function. Several distinct paradigms for loss function design have emerged.
Contrastive Learning Approaches: One major family of loss functions employs contrastive learning principles, directly shaping the embedding space by optimizing relative distances between samples. The seminal FaceNet [schroff2015facenet] introduces the triplet loss, designed to ensure that an anchor sample’s embedding is closer to its positive (same identity) counterpart than to any negative (different identity) sample by a predefined margin, typically in the Euclidean space.
While effective, triplet loss faces challenges in sampling informative triplets. Many randomly chosen triplets provide weak gradient signal, making training inefficient or necessitating complex hard-negative mining strategies. To address this, proxy-based methods are proposed. Techniques like Proxy Anchor Loss [kim2020proxy], originating from general deep metric learning, associate each class with learnable proxies (representative vectors), simplifying the loss computation by comparing samples to these proxies rather than exhaustively searching for pairs or triplets within a batch.
@@
DeepID2+ [sun2015deeply] 2014 Joint ID + verification loss for robust features Complex training with dual supervision losses
@@
CenterFace [centerloss] 2016 Center loss enhances intra-class compactness No explicit inter-class separation; needs tuning
@@
SphereFace [liu2017sphereface] 2017 Angular margin enforces hyperspherical separation Training instability from angular multiplicity
@@
L2-Face [ranjan2017l2] 2017 L2 norm constraint improves angular discrimination Needs careful radius tuning; no margin enforcement
@@
ArcFace [deng2019arcface] 2018 Additive angular margin boosts inter-class separation Fixed margin may hurt low-quality samples
@@
CosFace [wang2018cosface] 2018 Cosine margin improves class separability stably Uniform margin not adaptive; needs tuning
@@
SeqFace [hu2018seqface] 2018 Sequence-aware loss improves temporal supervision Needs sequence data; dual loss increases complexity
@@
Git Loss [calefati2018git] 2018 Unified softmax + center loss boosts discrimination Extra tuning and complexity with marginal gain
@@
MagFace [meng2021magface] 2021 Feature norm models quality for adaptive margin Complex loss and quality-norm assumptions
@@
AdaFace [kim2022adaface] 2022 Dynamic margin based on feature norm quality Relies on norm-quality link and tuning
@@
ElasticFace [boutros2022elasticface] 2022 Elastic margin adapts to feature variability Stochastic margins add tuning and training cost
@@
UniFace [zhou2023uniface] 2023 Similarity threshold improves verification alignment Global constraints increase optimization cost
@@
UniTSFace [jia2023unitsface] 2023 Sample-to-sample loss optimizes verification Pairwise loss and threshold learning cause overhead
@@
UCFace [ahn2024uncertainty] 2024 Uncertainty and probability density aware contrastive learning Cannot be used by itself, must be accompanied by margin loss
@@
LAFS [sun2024lafs] 2024 Landmark based SSL pretraining helps FR Loss depends on pretrained model and the landmark quality.
@@
Further refining the contrastive approach, Supervised Contrastive Learning (SupCon) [khosla2020supervised] generalizes the loss to leverage all positive samples available for an anchor within a batch, contrasting them against all negative samples. This more data-efficient approach has been successfully applied to FR, for instance, in UCFace [ahn2024uncertainty]. Other works adapt contrastive ideas for specific goals: Open-Set Biometrics [su2024open] focuses on improving open-set performance by explicitly minimizing scores between non-mated pairs, while CAFace [kim2022cluster] uses a contrastive-style cosine similarity loss to enforce consistency between embeddings of low-quality images and their high-quality counterparts, promoting quality invariance. Related efforts also explore optimizing embedding spaces to better align with recognition objectives, such as in [tu2009optimizing], where features are learned to directly improve performance metrics for identification and verification separately.
@@
Margin-based Softmax Losses: A dominant and highly successful approach in deep FR involves modifying the standard softmax cross-entropy loss to directly enhance feature discriminability. The core motivation is to learn embeddings that exhibit smaller intra-class variations (same person are close together) while simultaneously maximizing inter-class separation (different people are far apart).
@@
$\mathcal{L}_{CE}(\mathbf{x}_{i})=-\log\tfrac{\exp(\mathbf{W}_{y_{i}}^{\top}\mathbf{z}_{i}+b_{y_{i}})}{\sum_{j=1}^{C}\exp(\mathbf{W}_{j}^{\top}\mathbf{z}_{j}+b_{j})},$ (1)
@@
where $\mathbf{W}_{j}$ is the weight vector for the $j$-th class, $b_{j}$ is the bias term, and $C$ is the total number of classes or identities in the training set. While effective for classification, this formulation doesn’t explicitly enforce the metric learning objective crucial for FR where we encounter identities not seen during training.
@@
An early work moving in this direction is Center Loss [centerloss], which adds an auxiliary loss term to the standard softmax. This term penalizes the Euclidean distances between the deep features and their corresponding learned class centers, directly encouraging intra-class compactness. A significant breakthrough comes with the normalization of both feature embeddings ($\|\mathbf{z}_{i}\|=1$) and classification weights ($\|\mathbf{W}_{j}\|=1$, and setting $b_{j}=0$). This reformulation, pioneered by SphereFace [liu2017sphereface], maps the optimization problem onto a hypersphere where the dot product $\mathbf{W}_{j}^{T}\mathbf{z}_{i}$ becomes equivalent to $\cos\theta_{j}$, the cosine of the angle between the feature vector $\mathbf{z}_{i}$ and the weight vector $\mathbf{W}_{j}$. A scaling factor $s$ is typically introduced to control the radius of the hyperspherical feature space. The loss then becomes:
$\mathcal{L}_{cos}(\mathbf{x}_{i})=-\log\tfrac{\exp(s\!\cdot\!\cos\theta_{y_{i}})}{\sum_{j=1}^{C}\exp(s\!\cdot\!\cos\theta_{j})}.$ (2)
@@
Building on this normalized angular space, the key innovation is the introduction of explicit margins to make the learning objective more stringent. CosFace [wang2018cosface] introduces an additive cosine margin ($m$) by modifying the target logit to $s\cdot(\cos\theta_{y_{i}}-m)$. ArcFace [deng2019arcface] proposes an additive angular margin ($m$) by modifying the target angle itself, resulting in a target logit of $s\cdot\cos(\theta_{y_{i}}+m)$. Both approaches effectively create a decision boundary gap, forcing learned features for the same identity to cluster more tightly in the angular space, thereby significantly improving discriminative power. A visual comparison of contrastive triplet loss and margin-based normalized softmax loss is illustrated in Fig. 8, highlighting how each paradigm optimizes the embedding space to enhance FR performance.
@@
Subsequent research further refines these margin-based concepts. MagFace [meng2021magface] proposes leveraging the magnitude of the feature vector (before normalization) as an indicator of face image quality, incorporating an auxiliary loss to promote larger magnitudes for higher-quality samples. AdaFace [kim2022adaface] addresses the challenge posed by low-quality or difficult samples by introducing an adaptive margin function. It dynamically adjusts the margin stringency based on image quality indicators, reducing the negative impact of potentially unrecognizable faces in the training process.
@@
These advancements in margin-based softmax losses lead to remarkable performance gains, pushing verification accuracy on high-quality benchmarks like LFW [lfw] and CFP-FP [cfpfp] towards saturation (often exceeding 99%). This success shifts the community’s focus towards improving performance in more challenging, unconstrained scenarios, particularly those involving low-quality images, as represented by benchmarks like IJB-S,TinyFace or BRIAR [ijbs, cheng2018low, BRIAR]. The IJB-S and BRIAR datasets are limited-access evaluation sets developed under U.S. Government research programs and are distributed selectively for research use.
@@
Further refinements continued to explore margin dynamics; for example, ElasticFace [boutros2022elasticface] introduces randomized margins for greater flexibility during training, while UniFace [zhou2023uniface] proposes the Unified Cross-Entropy (UCE) loss specifically aiming to guarantee a clear separation threshold between positive and negative pairs.
@@
Margin-based softmax variants [deng2019arcface, kim2022adaface, zhou2023uniface] currently dominate SoTA results. Contrastive methods remain a helpful auxiliary loss, on top of margin-based softmax losses. A summary of various loss functions are shown in Tab. I.
@@
A significant effort has focused on improving model interpretability, e.g., understand how the network makes decisions. Towards this, Yin et al. [yin2019towards] propose spatial and feature activation diversity losses. These encourage the network to learn more structured representations where different spatial activations may correspond to different facial aspects, while also making these interpretable features discriminative and robust to occlusions. Similarly, the Explainable Channel Loss (ECLoss), also framed as Activation Template Matching Loss [Lin2024ECLoss], encourages specific channels within convolutional layers to specialize in detecting distinct facial parts (e.g., eyes) without explicit part annotations, thereby providing a direct interpretation of channel function.
Knowledge distillation (KD) offers another avenue, often targeting specific challenges. For low-resolution FR, Attention Similarity KD (A-SKD) [shin2022teaching] transfers teacher attention maps to guide the student’s focus. For efficient large-scale training, Li et al. [li2023rethinking] develop feature-based KD techniques, like reverse distillation, that importantly remove the need for identity supervision for the student, saving memory while addressing the teacher-student ’intrinsic gap’. These examples highlight the use of specialized loss functions and KD strategies to imbue models with desirable properties like explainability, robustness, or training efficiency, complementing the primary task of recognition.
Figure 9: Plots showing the growth of FR datasets over time. The left illustrates the number of images (in millions), and the right shows the number of subjects (in thousands) for each dataset.
@@ 4.2 Datasets
Before deep learning, FR was limited by small datasets. Prior to the 1996 FERET database, which had 1,199 individuals and became a major benchmark [phillips1998feret, phillips2000feret], most studies used fewer than 50 subjects, restricting generalization under unconstrained conditions. These small datasets sufficed for classical approaches like PCA based Eigenfaces [turk1991face], which relied on analytical solutions. The advent of deep learning [krizhevsky2012imagenet, taigman2014deepface, schroff2015facenet] created a vast need for data, as neural networks require far more samples to learn model parameters [vgg, ms1m, webface].
The availability and scale of training data have been pivotal factors driving the remarkable progress in deep learning [krizhevsky2012imagenet]. Publicly available large-scale face datasets (e.g., MS-Celeb-1M, VGGFace2) spurred rapid advancements in the mid-2010s. Fig. 10 provides an overview of several influential datasets commonly used in the field, detailing the number of images and unique identities they contain. A clear trend emerges from this summary: a dramatic increase in dataset size over time. Early benchmark datasets like CASIA-WebFace [casia] offer around half a million images from ten thousand subjects. In contrast, subsequent collections such as MS1MV2/V3 [ms1m], Glint360K [an2021partial], and particularly the WebFace series [webface], have pushed these numbers significantly higher, culminating in WebFace42M with over 42 million face images spanning 2 million identities. This growth reflects the community’s understanding that larger and more diverse datasets are crucial for training accurate and generalizable FR models. It is important to note that MS-Celeb-1M and VGGFace2 have been discontinued by the dataset creators. Some popular public datasets and pretrained model checkpoints are available in this link.
@@
It is important to note the origin and labeling methodology of many of these large-scale datasets. A significant portion, including prominent datasets like MS-Celeb-1M and the WebFace series, are curated by collecting images from publicly accessible sources on the Internet, often leveraging search engines or social media platforms. Consequently, the identity labels associated with these images are frequently “pseudo-labels,” because web searches of celebrities may return different subject images. Due to the volumn of these datasets, the labels are generated through automated clustering algorithms or matching, rather than manual verification. While efforts are made to clean and refine these labels, noise and inaccuracies can persist. Some approaches, like that used for WebFace260M [webface], employ iterative self-labeling and retraining of specialized labeler models to improve the quality of these pseudo-labels over multiple cycles. Since benchmark datasets [lfw, cfpfp, agedb, ijbb, ijbc] are also curated from public web, training datsets need to ensure that the identities in training and test sets do not overlap.
@@
Also, the practice of web-scraping facial images raises significant privacy and ethical concerns within the research community and society at large [wang2024beyond], as individuals may not have provided explicit consent for their images to be used in this manner for developing and training recognition systems. This issue remains an active area of discussion and necessitates careful consideration of data governance and ethical guidelines moving forward.
@@
In addition to the large-scale 2D image datasets collected primarily from the web, the field also utilizes 3D face datasets [blanz1999morphable, paysan2009bfm, gerig2018morphable, booth2018large, ploumpis2019combining, abrevaya2018multilinear, albrecht2013statistical, li2017learning, ranjan2018generating, li2020learning, wang2022faceverse, schwartz2010inface]. These datasets capture the geometric structure of the face, often along with texture information, using specialized acquisition techniques like 3D scanners, structured light, or multi-camera stereo systems. A key distinction is that 3D datasets are typically collected under controlled laboratory settings with the explicit consent of the participants. This controlled acquisition allows for high-quality, precise capture of facial shape in RGBd (depth), which can offer inherent robustness advantages against variations in pose and illumination compared to 2D images.
@@
3D face recognition has also emerged as a parallel research track, with dedicated benchmarks and evaluation campaigns [phillips2005overview]. Notably, datasets such as FRGC [phillips2005overview] and Lock3DFace [zhang2016lock3dface], which leverages low-cost RGB-D sensors like Kinect, have been widely used in the community to advance recognition algorithms under realistic conditions.
@@
However, the process of 3D data acquisition is significantly more complex, time-consuming, and expensive. Consequently, the volume of available 3D face data, both in terms of the number of scans and the number of unique subjects, is substantially smaller compared to the massive scale achieved by web-scraped 2D datasets. This difference in scale limits the use of 3D face datasets for training the deep models that dominate current FR research, although they remain valuable for specific research tasks, evaluation, and applications where 3D information is critical. For general FR applications, such as in law enforcement, immigration, or airport screening, RGB cameras offer a more practical solution considering legacy databases and return on investment.